# DeformTime: Capturing Variable Dependencies with Deformable Attention for Time Series Forecasting

### v2

@@ Abstract
In multivariate time series (MTS) forecasting, existing state-of-the-art deep learning approaches tend to focus on autoregressive formulations and overlook the information within exogenous indicators. To address this limitation, we present DeformTime, [^1] The source code of DeformTime is available at github.com/ClaudiaShu/DeformTime. a neural network architecture that attempts to capture correlated temporal patterns from the input space, and hence, improve forecasting accuracy. It deploys two core operations performed by deformable attention blocks (DABs): learning dependencies across variables from different time steps (variable DAB), and preserving temporal dependencies in data from previous time steps (temporal DAB). Input data transformation is explicitly designed to enhance learning from the deformed series of information while passing through a DAB. We conduct extensive experiments on $6$ MTS data sets, using previously established benchmarks as well as challenging infectious disease modelling tasks with more exogenous variables. The results demonstrate that DeformTime improves accuracy against previous competitive methods across the vast majority of MTS forecasting tasks, reducing the mean absolute error by $10\%$ on average. Notably, performance gains remain consistent across longer forecasting horizons.
@@ 1 Introduction
Time series forecasting models provide multifaceted solutions for many application domains, including health [39, 15], climate [37], energy [7], transport [24], and finance [38]. In their pursuit for greater accuracy, forecasting methods have always been trying to benefit from multimodal exogenous predictors [6, 1]. Over recent years, a plethora of new digitised information sources has become available, accompanied by a rapid development of capable and efficient deep learning architectures [35, 19]. However, current state-of-the-art (SOTA) models for multivariate time series (MTS) forecasting favour the deployment of many-to-many sequence deep learning architectures [32, 46, 29, 44]. We argue that this modelling approach does not necessarily provide the right learning mechanism for tasks with a definitive target (endogenous) variable and a considerable amount of exogenous variables. Furthermore, although various neural networks (NNs) were proposed to address MTS forecasting challenges [55, 51, 13, 29], many [53, 32, 16, 23, 17] were not designed to incorporate inter-variate dependencies, a key property of conventional forecasters [34, 14, 3].
@@
There certainly exist neural MTS forecasting models that attempt to leverage dependencies across variables and time. Solutions based on recurrent neural network (RNN) architectures [19, 25] can capture information from previous time steps (up to an extent), but have a limited capacity in establishing inter-variate dependencies. LightTS [54] uses multi-layer perceptrons to model dependencies within the input data, first along the temporal dimension and then across variables. For transformer-based models, Crossformer [55] proposes to first partition the input into temporal patches, and then use a router module to capture information across variables from different time steps, improving performance compared to pre-established RNN and Graph NN architectures [19, 47]. However, both LightTS and Crossformer have been outperformed by later proposed models that do not establish dependencies between variables [53, 32, 16, 23, 17]. The counterargument from these approaches is that forecasters without inter-variable dependency modules benefit from being able to use longer look-back windows without significantly increasing model complexity [11]. Nevertheless, a common issue throughout the deep learning MTS forecasting literature that casts doubt on some of these conclusions is the inappropriateness of various benchmark tasks (see Appendix A.5). In addition, recent work has argued that to improve performance while exploiting inter-variable dependencies, more effort is required in temporally aligning the input time series [56]. Motivated by this, we have introduced guided re-arrangements of the input to better capture inter- and intra-variable dynamics.
@@
Deformable neural networks were initially proposed in computer vision [5] to accommodate geometric transformations with Convolutional Neural Networks (CNNs). Combined with transformer-based NN architectures, deformation has achieved SOTA performance in various tasks [60, 4, 48]. Combining our aforementioned insights with the concept of deformation has led to the development of DeformTime, a novel MTS forecasting model that deploys a deformable module on top of transformer encoders to introduce some flexibility in the determination of receptive fields across different variables and time steps. The premise of DeformTime is the inclusion of learnable mechanisms that we refer to as deformable attention blocks (DABs). These enable transformations of the input information stream that enhance learning from key patterns within endo- but most importantly exogenous variables, an operation that ultimately improves forecasting accuracy.
@@
    We propose DeformTime, a novel MTS forecasting model, that better captures inter- and intra-variate dependencies at different temporal granularities. It comprises two DABs which facilitate learning from adaptively transformed input across variables (V-DAB) and time (T-DAB).
@@
    We assess the predictive accuracy of MTS forecasters on $3$ established benchmarks as well as $3$ new infectious disease prevalence tasks, each across $4$ increasing horizons. We argue that the disease prevalence tasks support a more thorough evaluation because data sets include a substantial amount of exogenous variables, and encompass multiple years, $4$ of which are used as distinct annual test periods.
@@
    In our experiments, DeformTime reduces the mean absolute error (MAE) by $10\%$ on average compared to the most competitive forecasting method (can be a different one each time). There is a $7.4\%$ MAE reduction based on established benchmarks and $12.6\%$ for the disease forecasting tasks. Overall, performance gains remain stable as the forecasting horizon increases.
@@ 2 MTS forecasting task definition
Across our experiments, we consider an MTS forecasting task where the focus is on a single target variable, i.e. there might exist multiple inputs, and optionally multiple outputs, but we are only concerned with the predictions of one output variable. All models follow a rolling window forecasting setting with a look-back window of $L$ time steps. To be more precise, at time step $t$, $\mathbf{Q}_{t}\!\in\!\mathbb{R}^{L\times C}$ holds the time series of $C$ exogenous variables over the $L$ time steps $\{t\!-\!L\!+\!1,\dots,t\!-\!1,t\}$. In addition, $\mathbf{y}_{t-\delta}\!\in\!\mathbb{R}^{L}$ holds the corresponding $L$ autoregressive historical values for the target variable for time steps $\{t\!-\!\delta\!-\!L\!+\!1,\dots,t\!-\!\delta\!-\!1,t\!-\!\delta\}$. Note that $\delta\!\in\!\mathbb{N}_{0}$ introduces an optional delay of $\delta$ time steps between the observed exogenous variables and the endogenous (target) variable. This becomes relevant in a real-time infectious disease forecasting task, where estimates for the rate of an infectious disease are becoming available with a delay of $\delta=7$ or $14$ days, when other indicators are not affected by this (see Appendix A.2). Joining both input signals, we obtain $\mathbf{Z}_{t}=\left[\mathbf{Q}_{t},\mathbf{y}_{t-\delta}\right]\!\in\!\mathbb{R}^{L\times\left(C+1\right)}$. Commonly, the prediction target is denoted by $\mathbf{y}_{t+H}\!\in\!\mathbb{R}^{H}$, where $H$ indicates the number of time steps we are forecasting ahead. However, some (baseline) models follow a multi-task learning formulation whereby every input variable becomes a prediction target. On this occasion, the prediction targets are denoted by $\mathbf{Y}_{t+H}=\left[\mathbf{Q}_{t+H},\mathbf{y}_{t+H}\right]\!\in\!\mathbb{R}^{H\times\left(C+1\right)}$. Hence, the aim of the forecasting task is to learn a function $f\!\left(\cdot\right)$ such that $f\!:\mathbf{Z}_{t}\rightarrow\mathbf{y}_{t+H}$ or $\mathbf{Y}_{t+H}$. Irrespective of the number of outputs, forecasting accuracy is reported on the final estimate of $\mathbf{y}_{t+H}$, $y_{t+H}\!\in\!\mathbb{R}$, which denotes the value of the target variable at time step $t\!+\!H$. For notational simplicity, we have chosen to only imply temporal subscripts for the remainder of the manuscript (i.e. use $\mathbf{Z}$ over $\mathbf{Z}_{t}$).
@@ 3 Time series forecasting with DeformTime
An overview of DeformTime’s structure is presented in Figure 1. Motivated by our assumption that variables (endo- and exogenous) in an MTS task may not only be correlated with the historical values of the target variable, but also with other predictors within adjacent time steps, DeformTime deploys a modified transformer encoder layer with two core modules that attempt to account for both dependencies: a Variable Deformable Attention Block (V-DAB) to capture inter-variable dependencies across time, and a Temporal Deformable Attention Block (T-DAB) to capture intra-variable dependencies across different time periods. In the following sections, we provide a detailed description of all modules and operations of DeformTime.
@@ 3.1 Multi-head attention
We first revisit the multi-head attention mechanism originally proposed by Vaswani et al. [41]. Given an input $\mathbf{Z}\!\in\!\mathbb{R}^{v\times u}$, $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}\!\in\!\mathbb{R}^{v\times u}$ respectively denote the query, key, and value embeddings projected from $\mathbf{Z}$ with learnable linear projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K},$ and $\mathbf{W}_{V}\!\in\!\mathbb{R}^{u\times u}$ (as in $\mathbf{Q}=\mathbf{Z}\mathbf{W}_{Q}$). Assuming $M$ attention heads, these embeddings are each partitioned into $M$ non-overlapping submatrices column-wise. For example, the $m$-th submatrix of $\mathbf{Q}$, denoted by $\mathbf{Q}_{m}$, has a dimensionality of $v\times(u/M)$. Following up on this notational convention, the output of the $m$-th attention head, $\mathbf{A}_{m}\!\in\!\mathbb{R}^{v\times(u/M)}$, is given by
@@
3.2 Neighbourhood-aware input embedding
The receptive field of DeformTime is determined by a CNN with sampling offset (see section 3.3). This introduces a limitation in capturing relationships between neighbouring variables. We seek to minimise the impact of this by learning a Neighbourhood-Aware input Embedding (NAE) as follows. We first re-arrange the order of variables in the input $\mathbf{Z}$, ranking them based on their linear correlation with the target variable in a temporally aligned fashion (see also Appendix C.1). We then embed the correlation-driven re-arranged version of the input that captures $C+1$ variables to a hidden dimension $d$. However, we do not perform this embedding as a single holistic step, but embed $G$ neighbouring groups of variables, each one to a shortened embedding of size $d/G$, using a fully connected layer. [^2] The input is zero-padded to a proper length, if the number of variables is not divisible by $G$. Both $d$ and $G$ are hyperparameters that we learn during training. The embeddings of each input variable grouping are concatenated to form $\mathbf{E}\!\in\!\mathbb{R}^{L\times d}$. Finally, we also add a sinusoidal position embedding [41, 57], $\mathbf{P}_{\texttt{n}}\!\in\!\mathbb{R}^{L\times d}$, to $\mathbf{E}$, and normalise within layer as widely adopted in prior work [8, 55]. Hence, the embedding $\mathbf{Z}_{\texttt{e}}\!\in\!\mathbb{R}^{L\times d}$ of the input $\mathbf{Z}$ is given by
@@ 3.2 Neighbourhood-aware input embedding
where $\texttt{LN}(\cdot)$ denotes layer normalisation [2].
@@ 3.3 Variable deformable attention block (V-DAB)
To determine the offset $\Delta\mathbf{p}\!\in\!\mathbb{R}^{2}$ we conduct the following series of operations. We first obtain a query embedding, $\mathbf{Q}_{\texttt{p}}\!\in\!\mathbb{R}^{\ell\times d}$, over a patch of the input as described in section 3.1, i.e. $\mathbf{Q}_{\texttt{p}}=\mathbf{Z}_{\texttt{p}}\mathbf{W}_{Q}$. We then pass $\mathbf{Q}_{\texttt{p}}$ from a 2D CNN, denoted by $\theta_{\texttt{off}}$, comprising a convolutional layer with a $k\!\times\!k$ kernel that captures neighbouring information, and a $1\!\times\!1$ convolutional layer that projects the embedded feature. We use the $\tanh$ activation function and a learnable amplitude $\alpha\!\in\!\mathbb{R}_{>0}$ (that controls the upper bound of the sampling range) to obtain $\Delta\mathbf{p}$. Hence, $\Delta\mathbf{p}$ is given by
@@
We use $\Delta\mathbf{p}$ and bilinear interpolation ($\psi$) over a $2\!\times\!2$ grid determined by $\mathbf{p}+\Delta\mathbf{p}$ [5], to obtain the deformed patch $\mathbf{Z}_{\texttt{d}}$:
@@
where $\mathbf{P}_{\texttt{v}}\!\in\!\mathbb{R}^{\ell\times d}$ is a relative positional bias [40] that is added to assist in maintaining some pairwise positional information after the deformation.
@@
where $\mathbf{W}_{i}\!\in\!\mathbb{R}^{d\times d}$ is the weight matrix, to form the attention embedding $\mathbf{A}_{i}\!\in\!\mathbb{R}^{\ell\times d}$ for the $i$-th segmented patch. The output from all patches is concatenated across the temporal dimension, forming $\mathbf{A}\!\in\!\mathbb{R}^{(n\cdot\ell)\times d}$. The final output of the V-DAB module, $\mathbf{Z}_{\texttt{v}}\!\in\!\mathbb{R}^{L\times d}$, is given by

### v3

@@ Abstract
In multivariable time series (MTS) forecasting, existing state-of-the-art deep learning approaches tend to focus on autoregressive formulations and often overlook the potential of using exogenous variables in enhancing the prediction of the target endogenous variable. To address this limitation, we present DeformTime, a neural network architecture that attempts to capture correlated temporal patterns from the input space, and hence, improve forecasting accuracy. It deploys two core operations performed by deformable attention blocks (DABs): learning dependencies across variables from different time steps (variable DAB), and preserving temporal dependencies in data from previous time steps (temporal DAB). Input data transformation is explicitly designed to enhance learning from the deformed series of information while passing through a DAB. We conduct extensive experiments on $6$ MTS data sets, using previously established benchmarks as well as challenging infectious disease modelling tasks with more exogenous variables. The results demonstrate that DeformTime improves accuracy against previous competitive methods across the vast majority of MTS forecasting tasks, reducing the mean absolute error by $7.2\%$ on average. Notably, performance gains remain consistent across longer forecasting horizons.
@@ 1 Introduction
Time series forecasting models provide multifaceted solutions for many application domains, including health (Shaman & Karspeck, 2012; Ioannidis et al., 2022), climate (Reichstein et al., 2019), energy (Deb et al., 2017), transport (Li et al., 2018), and finance (Sezer et al., 2020). In their pursuit for greater accuracy, forecasting methods have always been trying to benefit from multimodal exogenous predictors (De Gooijer & Hyndman, 2006; Athanasopoulos et al., 2011). Over recent years, a plethora of new digitised information sources has become available, accompanied by a rapid development of capable and efficient deep learning architectures (Rangapuram et al., 2018; Lai et al., 2018). Current state-of-the-art (SOTA) time series forecasting models are drawing particular focus on the deployment of many-to-many problem formulations (Nie et al., 2023; Wu et al., 2023; Luo & Wang, 2024b; Wang et al., 2024a). However, this approach does not necessarily provide the right learning mechanism for multivariable time series (MTS) forecasting (Hidalgo & Goodman, 2013), where we have a definitive target (endogenous) variable and a considerable amount of external (exogenous) input variables. Furthermore, although various neural network (NN) architectures were proposed to address forecasting challenges with multiple inputs (Zhang & Yan, 2023; Yi et al., 2023a; Huang et al., 2023; Luo & Wang, 2024b), many (Zeng et al., 2023; Nie et al., 2023; Jia et al., 2023; Lee et al., 2024; Jin et al., 2024; Lin et al., 2024) were not designed to incorporate inter-variable dependencies, a key property of conventional forecasting models (Pi & Peterson, 1994; Hyndman & Ullah, 2007; Bontemps et al., 2008).
@@
There certainly exist neural forecasting models that attempt to leverage dependencies across variables and time. Solutions based on recurrent neural network (RNN) architectures (Lai et al., 2018; Lin et al., 2023) can capture information from previous time steps (up to an extent), but have a limited capacity in establishing inter-variable dependencies. LightTS (Zhang et al., 2022) uses multi-layer perceptrons to model dependencies within the input data, first along the temporal dimension and then across variables. For transformer-based models, Crossformer (Zhang & Yan, 2023) proposes to first partition the input into temporal patches, and then use a router module to capture information across variables from different time steps, improving performance compared to pre-established RNN and Graph NN architectures (Lai et al., 2018; Wu et al., 2020). However, both LightTS and Crossformer have been outperformed by later proposed models that do not establish dependencies between variables (Zeng et al., 2023; Nie et al., 2023; Jia et al., 2023; Lee et al., 2024; Jin et al., 2024; Lin et al., 2024). The counterargument from these approaches is that models without inter-variable dependency modules benefit from being able to use longer look-back windows without significantly increasing model complexity (Han et al., 2023). Nevertheless, a common issue throughout the deep learning forecasting literature that casts doubt on some of these conclusions is the inappropriateness of various benchmark tasks (see Appendix A.4). In addition, recent work has argued that to improve performance while exploiting inter-variable dependencies, more effort is required in temporally aligning the input time series (Zhao & Shen, 2024). Attempts have also been made to capture dependencies with large receptive fields within segmented time series patches (Luo & Wang, 2024b). Motivated by this, we have introduced guided re-arrangements of the input to better capture inter- and intra-variable dynamics.
@@
Deformable neural networks were initially proposed in computer vision (Dai et al., 2017) to accommodate geometric transformations with Convolutional Neural Networks (CNNs). Combined with transformer-based NN architectures, deformation has achieved SOTA performance in various tasks (Zhu et al., 2021; Chen et al., 2021; Xia et al., 2022). Prior work (Wang et al., 2024c) has also deployed deformable mechanisms within time series forecasting. However, their application was limited to establishing intra-variable dependencies, a structural design also shared in (Luo & Wang, 2024a), making them less effective in capturing correlations between multiple variables over time. Importantly, this model was outperformed by later ones (Nie et al., 2023; Liu et al., 2024). Based on the aforementioned insights, this paper introduces DeformTime, a novel MTS forecasting model that deploys a deformable module on top of transformer encoders to introduce some flexibility in the determination of receptive fields across different variables and time steps. The premise of DeformTime is the inclusion of learnable mechanisms that we refer to as deformable attention blocks (DABs). These enable transformations of the input information stream that enhance learning from key patterns within endo- but most importantly exogenous variables, an operation that ultimately improves forecasting accuracy (for the target endogenous variable). Additionally, input is transformed in alignment with the operational goal of the DAB modules, guiding the modelling of dependencies within time intervals and at various temporal resolutions. Our model also deploys both fixed and learnable positional embeddings to preserve the sequential order of the input time series and capture relative positional information after deformation, respectively. These embeddings work in tandem within the proposed architecture.
@@
    We propose DeformTime, a novel MTS forecasting model, that better captures inter- and intra-variable dependencies at different temporal granularities. It comprises two DABs which facilitate learning from adaptively transformed input across variables (V-DAB) and time (T-DAB).
@@
    We assess the predictive accuracy of MTS forecasting models on $3$ established benchmarks as well as $3$ new infectious disease prevalence tasks, each across $4$ increasing horizons. We argue that the disease prevalence tasks support a more thorough evaluation because data sets include a substantial amount of exogenous variables, and encompass multiple years, $4$ of which are used as distinct annual test periods.
@@
    In our experiments, DeformTime reduces the mean absolute error (MAE) by $7.2\%$ on average compared to the most competitive forecasting method (which may be a different one for each evaluation). There is a $5\%$ MAE reduction based on established benchmarks and $9.3\%$ for the disease forecasting tasks. Overall, performance gains remain stable as the forecasting horizon increases.
@@ 2 MTS forecasting task definition
Across our experiments, we consider an MTS forecasting task where the focus is on a single target variable, i.e. there might exist multiple inputs, and optionally multiple outputs, but we are only concerned with the predictions of one output variable. All models follow a rolling window forecasting setting with a look-back window of $L$ time steps. To be more precise, at time step $t$, $\mathbf{Q}_{t}\!\in\!\mathbb{R}^{L\times C}$ holds the time series of $C$ exogenous variables over the $L$ time steps $\{t\!-\!L\!+\!1,\dots,t\!-\!1,t\}$. In addition, $\mathbf{y}_{t-\delta}\!\in\!\mathbb{R}^{L}$ holds the corresponding $L$ autoregressive historical values for the target variable for time steps $\{t\!-\!\delta\!-\!L\!+\!1,\dots,t\!-\!\delta\!-\!1,t\!-\!\delta\}$. Note that $\delta\!\in\!\mathbb{N}_{0}$ introduces an optional delay of $\delta$ time steps between the observed exogenous variables and the endogenous (target) variable. This becomes relevant in a real-time infectious disease forecasting task, where estimates for the rate of an infectious disease are becoming available with a delay of $\delta=7$ or $14$ days, when other indicators are not affected by this (see Appendix A.2). Joining both input signals, i.e. the exogenous and endogenous (autoregressive) variables, we obtain the input matrix $\mathbf{Z}_{t}=\left[\mathbf{Q}_{t},\mathbf{y}_{t-\delta}\right]\!\in\!\mathbb{R}^{L\times\left(C+1\right)}$. Commonly, the prediction target is denoted by $\mathbf{y}_{t+H}\!\in\!\mathbb{R}^{H}$, where $H$ indicates the number of time steps we are forecasting ahead. However, some (baseline) models follow a multi-task learning formulation whereby every input variable becomes a prediction target. On this occasion, the prediction targets are denoted by $\mathbf{Y}_{t+H}=\left[\mathbf{Q}_{t+H},\mathbf{y}_{t+H}\right]\!\in\!\mathbb{R}^{H\times\left(C+1\right)}$. Hence, the aim of the forecasting task is to learn a function $f\!\left(\cdot\right)$ such that $f\!:\mathbf{Z}_{t}\rightarrow\mathbf{y}_{t+H}$ or $\mathbf{Y}_{t+H}$. Irrespective of the number of outputs, forecasting accuracy is reported on the final estimate of $\mathbf{y}_{t+H}$, $y_{t+H}\!\in\!\mathbb{R}$, which denotes the value of the target variable at time step $t\!+\!H$. For notational simplicity, we have chosen to only imply temporal subscripts for the remainder of the manuscript (i.e. use $\mathbf{Z}$ over $\mathbf{Z}_{t}$).
@@ 3 Time series forecasting with DeformTime
An overview of DeformTime’s structure is presented in Figure 1. Motivated by our assumption that variables (endo- and exogenous) in an MTS task may not only be correlated with the historical values of the target variable, but also with other predictors within adjacent time steps, DeformTime deploys a modified transformer encoder layer with two core modules that attempt to account for both dependencies: a Variable Deformable Attention Block (V-DAB) to capture inter-variable dependencies across time, and a Temporal Deformable Attention Block (T-DAB) to capture intra-variable dependencies across different time periods. In the following sections, we provide a detailed description of all modules and operations of DeformTime. [^1] The source code of DeformTime is available at github.com/ClaudiaShu/DeformTime.
@@ 3.1 Multi-head attention
We first revisit the multi-head attention mechanism originally proposed by Vaswani et al. (2017). Given an input $\mathbf{Z}\!\in\!\mathbb{R}^{v\times u}$, $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}\!\in\!\mathbb{R}^{v\times u}$ respectively denote the query, key, and value embeddings projected from $\mathbf{Z}$ with learnable linear projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K},$ and $\mathbf{W}_{V}\!\in\!\mathbb{R}^{u\times u}$ (as in $\mathbf{Q}=\mathbf{Z}\mathbf{W}_{Q}$). Assuming $M$ attention heads, these embeddings are each partitioned into $M$ non-overlapping submatrices column-wise. For example, the $m$-th submatrix of $\mathbf{Q}$, denoted by $\mathbf{Q}_{m}$, has a dimensionality of $v\times(u/M)$. Following up on this notational convention, the output of the $m$-th attention head, $\mathbf{A}_{m}\!\in\!\mathbb{R}^{v\times(u/M)}$, is given by
@@
3.2 Neighbourhood-aware input embedding (NAE)
The receptive field of DeformTime is determined by a CNN with sampling offset (see section 3.3). This introduces a limitation in capturing relationships between neighbouring variables. We seek to minimise the impact of this by learning a neighbourhood-aware input embedding (NAE) as follows. We first re-arrange the order of variables in the input $\mathbf{Z}$, ranking them based on their linear correlation with the target variable in a temporally aligned fashion (see also Appendix D.1). We then embed the correlation-driven re-arranged version of the input that captures $C+1$ variables to a hidden dimension $d$. However, we do not perform this embedding as a single holistic step, but embed $G$ neighbouring groups of variables, each one to a shortened embedding of size $d/G$, using a fully connected layer. [^2] The input is zero-padded to a proper length, if the number of variables is not divisible by $G$. Our expectation is that the presence of stronger intra-group correlations will be subsequently leveraged by our model. Both $d$ and $G$ are hyperparameters that we learn during training. The embeddings of each input variable grouping are concatenated to form $\mathbf{E}\!\in\!\mathbb{R}^{L\times d}$. Finally, we also add a sinusoidal position embedding (Vaswani et al., 2017; Zhou et al., 2021), $\mathbf{P}_{\texttt{n}}\!\in\!\mathbb{R}^{L\times d}$, to $\mathbf{E}$, and normalise within layer as widely adopted in prior work (Dosovitskiy et al., 2021; Zhang & Yan, 2023). Hence, the embedding $\mathbf{Z}_{\texttt{e}}\!\in\!\mathbb{R}^{L\times d}$ of the input $\mathbf{Z}$ is given by
@@ 3.2 Neighbourhood-aware input embedding (NAE)
where $\texttt{LN}(\cdot)$ denotes layer normalisation (Ba et al., 2016).
@@ 3.3 Variable deformable attention block (V-DAB)
To determine the position offset $\Delta\mathbf{p}\!\in\!\mathbb{R}^{2}$ we conduct the following series of operations. We first obtain a query embedding, $\mathbf{Q}_{\texttt{p}}\!\in\!\mathbb{R}^{\ell\times d}$, over a patch of the input as described in section 3.1, i.e. $\mathbf{Q}_{\texttt{p}}=\mathbf{Z}_{\texttt{p}}\mathbf{W}_{Q}$. We then pass $\mathbf{Q}_{\texttt{p}}$ from a 2D CNN, denoted by $\theta_{\texttt{off}}$, comprising a convolutional layer with a $k\!\times\!k$ kernel that captures neighbouring information, and a $1\!\times\!1$ convolutional layer that projects the embedded feature. We use the $\tanh$ activation function and a learnable amplitude $\alpha\!\in\!\mathbb{R}_{>0}$ (that controls the upper bound of the sampling range) to obtain $\Delta\mathbf{p}$. Hence, $\Delta\mathbf{p}$ is given by
@@
We use $\Delta\mathbf{p}$ and bilinear interpolation ($\psi$) over a $2\!\times\!2$ grid determined by $\mathbf{p}+\Delta\mathbf{p}$ (Dai et al., 2017), to obtain the deformed patch $\mathbf{Z}_{\texttt{d}}$:
@@
where $\mathbf{P}_{\texttt{v}}\!\in\!\mathbb{R}^{\ell\times d}$ is a relative positional bias (Shaw et al., 2018) that is added to assist in maintaining some pairwise positional information after the deformation.
@@
where $\mathbf{W}_{i}\!\in\!\mathbb{R}^{d\times d}$ is the weight matrix, to form the attention embedding $\mathbf{A}_{i}\!\in\!\mathbb{R}^{\ell\times d}$ for the $i$-th segmented patch. The output from all patches is concatenated across the temporal dimension, forming $\mathbf{A}\!\in\!\mathbb{R}^{(n\cdot\ell)\times d}$. We conduct cross-attention over the original latent variable captured in $\mathbf{Q}_{\texttt{p}}$ and the deformed latent variables $\mathbf{K}_{\texttt{d}},\mathbf{V}_{\texttt{d}}$. Specifically, elements at position $\mathbf{p}$ can attend to elements at position $\mathbf{p}+\Delta\mathbf{p}$. This enables the model to learn from interactions between neighbouring variables across time.
The final output of the V-DAB module, $\mathbf{Z}_{\texttt{v}}\!\in\!\mathbb{R}^{L\times d}$, is given by